OpenStudy (deadshot):

Divide 2x^2 + 17x + 35 by x + 5. A. x + 7 B. 2x + 7 C. x + 27 D. 2x + 27

OpenStudy (campbell_st):

so this is really \[\frac{(2x^2 + 17x + 35)}{(x + 5)}\] can you factor the numerator..?

OpenStudy (deadshot):

No, there is no GCF.

OpenStudy (campbell_st):

yes there is... do you know how to factorise a quadratic..?

OpenStudy (deadshot):

no

OpenStudy (hba):

I have a better idea for you, Use Synthetic division method.

OpenStudy (cwrw238):

the factors of this are of the form (2x + a )(x + b) where a * b = 35 and 2a + b = 17

OpenStudy (campbell_st):

why not just factorise \[\frac{(x + 5)(2x + 7)}{(x + 5)}\]

OpenStudy (cwrw238):

sorry 2b + a = 17

OpenStudy (cwrw238):

yes - its easier to just factorise if you are competent at this

OpenStudy (deadshot):

so then the (x + 5) in the denominator cancels out the (x + 5) in the numerator, leaving (2x + 7)?

OpenStudy (cwrw238):

right

OpenStudy (campbell_st):

thats correct... easy if you can factor..

OpenStudy (deadshot):

Okay, thanks!